### v1

# Does Privacy Always Harm Fairness? Data-Dependent Trade-offs via Chernoff Information Neural Estimation

Abstract
Fairness and privacy are two vital pillars of trustworthy machine learning. Despite extensive research on these individual topics, the relationship between fairness and privacy has received significantly less attention. In this paper, we utilize the information-theoretic measure Chernoff Information to highlight the data-dependent nature of the relationship among the triad of fairness, privacy, and accuracy. We first define Noisy Chernoff Difference, a tool that allows us to analyze the relationship among the triad simultaneously. We then show that for synthetic data, this value behaves in 3 distinct ways (depending on the distribution of the data). We highlight the data distributions involved in these cases and explore their fairness and privacy implications. Additionally, we show that Noisy Chernoff Difference acts as a proxy for the steepness of the fairness-accuracy curves. Finally, we propose a method for estimating Chernoff Information on data from unknown distributions and utilize this framework to examine the triad dynamic on real datasets. This work builds towards a unified understanding of the fairness-privacy-accuracy relationship and highlights its data-dependent nature.
1 Introduction
As machine learning (ML) systems are increasingly deployed in high-stakes settings, ensuring their trustworthiness has become critical. Among the many facets of trustworthy ML, fairness and privacy stand out as core concerns, yet both are frequently violated in practice. Unfairness has been documented across domains such as facial recognition (Garvie and Frankle, 2016), text-to-image generation (Friedrich et al., 2023; Bianchi et al., 2023), and predictive tasks like recidivism (Barocas and Selbst, 2016; Chouldechova, 2016) and loan approvals (Das et al., 2021), prompting a broad literature on fair learning (Hardt et al., 2016; Agarwal et al., 2018; Alghamdi et al., 2022). Privacy risks, highlighted by attacks that recover sensitive training data (Shokri et al., 2017; Carlini et al., 2021) and amplified by the rise of large-scale AI models (Gomstyn and Jonker, 2024), have likewise spurred extensive research on differentially private (DP) learning (Abadi et al., 2016; Papernot et al., 2018; Kang et al., 2020).
Yet while fairness and privacy have been deeply studied in isolation, their interaction remains less understood. Can models satisfy both simultaneously, or does enforcing one inevitably compromise the other? Recent work points to both incompatibility, exposing trade-offs between the two (Bagdasaryan et al., 2019; Chang and Shokri, 2021; Sanyal et al., 2022), and potential compatibility under specific assumptions (Cummings et al., 2019; Mangold et al., 2023).
In this work, we show that the answer ultimately lies in the structure of the data itself. We introduce the Noisy Chernoff Difference, a metric grounded in Chernoff Information, that captures the triad of fairness, privacy, and accuracy by quantifying how separability between groups evolves under privacy-preserving noise. Under Gaussian assumptions, we prove that fairness–privacy interactions fall into three distinct regimes: privacy may erode fairness, leave it unaffected, or even improve it, effectively granting “fairness for free” when distributional conditions align. To bridge theory with practice, we develop, to our knowledge, the first algorithm to estimate Chernoff Information directly from data, leveraging advances in density ratio estimation via neural networks (Choi et al., 2022), named Chernoff Information Neural Estimation (CINE in Section 5).
Beyond the Gaussian case, we validate our framework on three datasets, MNIST (LeCun et al., 1998), UCI Adult Income (Becker and Kohavi, 1996), and HSLS (Ingels et al., 2011), showing that the same distribution-driven patterns hold in practice. Our findings provide the first concrete evidence that whether fairness and privacy are in conflict or in harmony is not a universal law, but a property of the underlying data distribution. This reframes the fairness–privacy dilemma: it is not simply about balancing competing constraints, but about understanding when data itself enables both goals to be achieved in tandem, transforming how we think about building more prudent, trustworthy machine learning systems.
2 Related Works
Privacy and Fairness.
The relationship between fairness and privacy has attracted increasing attention in recent years. Bagdasaryan et al. (2019) show that differential privacy disproportionately harms the accuracy of underrepresented groups, while Chang and Shokri (2021) demonstrate that minority groups are more vulnerable to membership inference attacks. Cummings et al. (2019) prove the impossibility of achieving exact fairness and DP with nontrivial accuracy, but propose an algorithm that satisfies DP with approximate fairness under relaxed constraints. Sanyal et al. (2022) show that under long-tailed, imbalanced distributions and strict privacy, enforcing both fairness and privacy worsens accuracy, whereas Mangold et al. (2023) establish that fairness disparities between private and non-private models decay inversely with sample size. Several works also develop algorithms jointly ensuring fairness, privacy, and accuracy (Lyu et al., 2020; Lowy et al., 2023; Ghoukasian and Asoodeh, 2024; Jagielski et al., 2019). For a comprehensive survey, see Yao and Juarez (2025).
Chernoff Information.
Chernoff Information, introduced by Chernoff (1952), is closely tied to Bayes error bounds. Nielsen (2011, 2013, 2022) provides key properties we exploit throughout this work. Although information-theoretic tools have been widely applied to fairness and privacy (Ghassami et al., 2018), Chernoff Information itself has rarely been used in these domains. Dutta et al. (2020) show that disparities in Chernoff Information imply a trade-off between equal opportunity fairness and accuracy, while Ünsal and Önen (2024) define Chernoff Differential Privacy, where output distributions for neighboring inputs must have bounded Chernoff Information. To our knowledge, ours is the first work to connect Chernoff Information with the joint analysis of accuracy, fairness, and privacy, and to provide an estimator beyond synthetic settings.
3 Problem Setting and Background
3.1 Setting
Consider a binary classification setting in which our data is defined by continuous, non-sensitive features $X$, sensitive attributes $S=\{0,1\}$, and labels $Y=\{0,1\}$. Using these parameters, we can define our data distribution as a mixture of conditional distributions $P_{0}(x)=\Pr(X|S=0,Y=0)$, $P_{1}(x)=\Pr(X|S=0,Y=1)$, $Q_{0}(x)=\Pr(X|S=1,Y=0)$, and $Q_{1}(x)=\Pr(X|S=1,Y=1)$. From this definition, we can refer to our groups as $P$, to be the group where $S=0$ and $Q$, to be the group where $S=1$. Further, we make an equal prior assumption ($P(Y=0|S=s)=P(Y=1|S=s)$ for all $s$). We discuss a relaxation of this assumption in Section 6. For our model, we consider a split classifier setting where we train a different classifier for $P$ and $Q$. Under the assumption of infinite model complexity (and sufficient group information), any model will converge towards the split classifier setting (Wang et al., 2021).
Goal.
Our central aim is to examine how different underlying data distributions ($P_{0},P_{1},Q_{0},$ and $Q_{1}$) affect the fundamental trade-off among fairness, privacy, and accuracy for a classifier.
Privacy.
We consider differential privacy in this paper as defined below:
Algorithm 1 SGD with Gaussian Input Perturbation
1:Dataset $D$, batch size $m$, learning rate $\alpha$, noise variance $\eta^{2}$, initial parameters $\bm{\theta_{0}}$
2:Create noisy dataset $\widetilde{D}=\{\mathbf{\tilde{x}_{i}},y\}_{i=1}^{n}$ where $\mathbf{\tilde{x}_{i}}=\mathbf{x_{i}}+\bm{\xi_{i}}$ with $\bm{\xi_{i}}\sim\mathcal{N}(\bm{0},\eta^{2}\mathbf{I})$
3:for $t=1,2,\dots,T$ do
4:  Sample a mini-batch $\{\mathbf{x_{i}},y_{i}\}_{i=1}^{m}\sim\widetilde{D}$
5:  Compute stochastic gradient:
6:         $\mathbf{g_{t}}=\frac{1}{m}\sum_{i=1}^{m}\nabla_{\theta}\ell(\bm{\theta_{t-1}};\mathbf{\tilde{x}_{i}},y_{i})$
7:  Update parameters: $\bm{\theta_{t}}\leftarrow\bm{\theta_{t-1}}-\alpha\mathbf{g_{t}}$
8:end for
9:return $\theta_{T}$
Definition 1 (Differential Privacy).
A randomized learning algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-differentially private if, for every pair of neighboring datasets $D$ and $D^{\prime}$ that differ in at most one element, and for every measurable subset $S$ of the output space of $\mathcal{A}$,
$\Pr\!\bigl[\mathcal{A}(D)\in S\bigr]\;\leq\;e^{\varepsilon}\,\Pr\!\bigl[\mathcal{A}(D^{\prime})\in S\bigr]+\delta.$ (1)
Differentially private machine learning methods are commonly divided into three categories (Jarin and Eshete, 2022): (1) input perturbation, (2) gradient/objective perturbation, and (3) output perturbation. In this paper, we focus on input perturbation (Fukuchi et al., 2017; Kang et al., 2020), as it naturally aligns with our goal of examining the data-dependent interplay between fairness and privacy. Empirical results further suggest that input perturbation offers advantages in regimes requiring high privacy constraints (Zhao et al., 2020).
A standard approach to achieve $(\varepsilon,\delta)$-DP is the Gaussian mechanism (Dwork et al., 2014), which adds independent Gaussian noise $\sim\mathcal{N}(0,\eta^{2})$ to each coordinate. In Algorithm 1, we detail how the Gaussian mechanism is applied to the input perturbation framework. By a direct application of Theorem A.1 in Dwork et al. (2014), we obtain:
Lemma 1.
Assume $\|\mathbf{x}\|_{2}\leq 1$ for all $\mathbf{x}\in\mathbb{R}^{d}$. When $\eta^{2}\geq\frac{8\log(1.25/\delta)}{\varepsilon^{2}}$, Algorithm 1 is $(\varepsilon,\delta$)-DP with respect to feature-level neighboring datasets (Definition A.2).
This is a loose bound that translates guarantees from local DP to central DP (providing privacy at the feature level — See Appendix A). Recent work has refined the analysis (extending privacy to the more general notion of neighboring datasets [Definition A.4] and tightening this bound) under additional assumptions, such as strong convexity of the loss function (Kang et al., 2020) or a quadratic loss (Fukuchi et al., 2017). Nevertheless, the fundamental relationship between $\eta^{2}$ and $(\varepsilon,\delta)$ remains unchanged: larger $\eta^{2}$ corresponds to smaller $\varepsilon$ and $\delta$, indicating stronger privacy guarantees. We therefore use $\eta^{2}$ as our measure of privacy throughout the remainder of the paper.
Fairness.
Among many notions of fairness in machine learning (Barocas et al., 2023), a widely used notion is equal opportunity (EO) (Hardt et al., 2016; Zafar et al., 2017), which measures fairness as the difference in true positive rates (or false negative rates) between the groups ($P$ and $Q$)
$\text{EO}=|\text{TPR}(P)-\text{TPR}(Q)|.$ (2)
As Chernoff Information naturally connects to this notion of fairness, we will utilize this throughout this work to demonstrate how Chernoff Information can provide insights into traditional fairness-accuracy curves. The connection between Chernoff Information and EO is detailed in the following subsection.
3.2 Chernoff Difference (CD)
Here we introduce Chernoff Information and its relevant properties building towards Chernoff Difference and its noisy variant. For more discussion please refer to Appendix B. Recent work (Dutta et al., 2020) has demonstrated that Chernoff Information can characterize the fundamental trade-off between fairness and accuracy. Building on this, we will leverage Chernoff Information as a tool to quantify the relationship between fairness, privacy, and accuracy. We begin with the formal definition below.
Definition 2 (Chernoff Information (Chernoff, 1952)).
For two distributions $P_{0}(x)$ and $P_{1}(x)$ the Chernoff Information is given by:
$C(P_{0},P_{1})=-\inf_{u\in(0,1)}\log(\int P_{0}(x)^{1-u}P_{1}(x)^{u}dx).$ (3)
Chernoff Information (CI) can be interpreted as a divergence that quantifies the separability between $P_{0}$ and $P_{1}$ (Dutta et al., 2020; Nielsen, 2011). For large values of CI, two hypotheses $P_{0}$ and $P_{1}$ are more separable, which indicates easier classification. Conversely, smaller values of CI imply less separability between the two hypotheses, and thus a more difficult classification setting. This is demonstrated by the role of CI in bounding the error exponent of the Bayes optimal classifier.
Lemma 2 ((Nielsen, 2011)).
For hypotheses $P_{0}(x)$ under $Y=0$ and $P_{1}(x)$ under $Y=1$, CI bounds the error of the Bayes Optimal Classifier $H$:
$\Pr[H(X)\neq Y]\leq e^{-C(P_{0},P_{1})}.$
Thus, CI provides a tight characterization of how well the best possible classifier can perform given sufficiently large data (Appendix B). We next define Chernoff Difference (CD).
Definition 3 (Chernoff Difference).
The Chernoff Difference between group $P$ and group $Q$ is defined as follows:
$\text{CD}=\left|C(P_{0},P_{1})-C(Q_{0},Q_{1})\right|.$ (4)
CD serves as a metric that compares the classification difficulty between two groups. A smaller CD indicates that the groups have similar separability between their positive and negative classes, whereas a larger CD implies a large disparity in separability. A recent work (Dutta et al., 2020) used CD to show an impossibility result: if CD is not zero, it is impossible to achieve fairness without sacrificing accuracy. While Dutta et al. (2020) draws a connection between CD and fairness, their analysis was limited to whether CD equals zero or not. In this work, we go one step further by relating the value of CD to the steepness of the fairness–accuracy curves. Below we clarify this relation, how Chernoff Difference (CD) functions as a notion of fairness, and how it relates to Equal Opportunity (EO).
Remark 1 (Chernoff Difference as a Data Fairness Measure).
Unlike traditional fairness metrics, which quantify disparities induced by a particular trained classifier, CD is a data-level notion of fairness. The role of CD is not to evaluate a learned predictor, but to characterize the best achievable error behavior permitted by the data. It characterizes disparities that are inherent to the underlying data distributions $(P_{0},P_{1})$ and $(Q_{0},Q_{1})$, independent of any specific model. In this sense, CD measures data fairness, rather than model fairness. This perspective explains the relationship between CD and the Bayes-optimal classifier, and how the input distribution affects the fairness–privacy–accuracy triad. If CD is large, then even the Bayes-optimal classifier must exhibit substantially different error behavior across groups, implying that any attempt to equalize error impacts accuracy significantly.
This can be more formally examined by relating CD to error rates in the asymptotic regime. Formally, Chernoff Information equals the error exponent of the Bayes-optimal classifier,
$C(P_{0},P_{1})=\min\{E_{\mathrm{FP}}^{P},E_{\mathrm{FN}}^{P}\},$ (5)
where $E_{\mathrm{FP}}$ and $E_{\mathrm{FN}}$ are the false positive and false negative error exponents. Consequently, Chernoff Difference can be written as
$\mathrm{CD}=\bigl|\min\{E_{\mathrm{FP}}^{P},E_{\mathrm{FN}}^{P}\}-\min\{E_{\mathrm{FP}}^{Q},E_{\mathrm{FN}}^{Q}\}\bigr|.$ (6)
Using the asymptotic relationships $E_{FP}\approx-\frac{1}{n}\log\text{FPR}$ and $E_{FN}\approx-\frac{1}{n}\log\text{FNR}$, CD corresponds to a disparity in the dominant error rates on a log scale:
$\mathrm{CD}\approx\frac{1}{n}\log\left|\frac{\max\{\mathrm{FPR}(P),\mathrm{FNR}(P)\}}{\max\{\mathrm{FPR}(Q),\mathrm{FNR}(Q)\}}\right|.$ (7)
Thus, CD captures a geometric difference rather than the arithmetic difference used in common fairness metrics such as Equal Opportunity. If the maximum error for both $P$ and $Q$ occurs at the FNR, then CD effectively captures the best possible FNR gap achievable by any classifier given the underlying data. Thus, CD plays a role analogous to Equal Opportunity. Both quantify a FNR gap, but while Equal Opportunity measures an additive difference, CD measures a multiplicative one. While these notions are not equivalent, they are related in meaningful ways. In the case of a fixed classifier for a single group, the relationship between the additive and multiplicative notion of fairness is monotonic. In the more general case, CD can serve as a rough lower and upper bound for the arithmetic-difference fairness metric (Theorem B.1) which can be tightened with knowledge of the error rates. In the following section, we empirically demonstrate that CD provides insight into the behavior of the EO fairness accuracy curve.
Next, to incorporate our notion of privacy, we define a noisy variant of Chernoff Difference.
Definition 4 (Noisy Chernoff Difference).
For $\eta^{2}\geq 0$, we define the Noisy Chernoff Difference as:
$\widetilde{\text{CD}}_{\eta^{2}}=\left|C(\widetilde{P}_{0},\widetilde{P}_{1})-C(\widetilde{Q}_{0},\widetilde{Q}_{1})\right|.$ (8)
where $\widetilde{P}_{i}=P_{i}+\mathcal{N}(0,\eta^{2}\mathbf{I})$ and $\widetilde{Q}_{i}=Q_{i}+\mathcal{N}(0,\eta^{2}\mathbf{I})$
While CD provides insight into the relationship between fairness and accuracy, $\widetilde{CD}_{\eta^{2}}$ provides a single value that can capture the relationship between all three: fairness, accuracy, and privacy. More specifically, it allows us to analyze how adding noise (stronger privacy) affects both fairness and accuracy. We illustrate this relationship in the following section using Gaussian distributions.
4 Gaussian Example
While we later show how to estimate CD from various distributions, including synthetic Gaussian mixtures and real datasets, we first provide clean insights into how different input distributions affect the fairness–privacy–accuracy trade-off using simple isotropic Gaussian examples. We define the conditional data distributions as follows:
$P_{0}=\mathcal{N}(\mu_{0},\sigma^{2}\mathbf{I}),P_{1}=\mathcal{N}(\mu_{1},\sigma^{2}\mathbf{I})$ (9)
$Q_{0}=\mathcal{N}(\zeta_{0},\tau^{2}\mathbf{I}),Q_{1}=\mathcal{N}(\zeta_{1},\tau^{2}\mathbf{I}),$ (10)
This allows us to extend a result from Dutta et al. (2020) to derive a closed form expression for the Chernoff Difference.
Lemma 3.
When $P_{0}(x)\sim\mathcal{N}(\mu_{0},\sigma^{2}\mathbf{I})$, $P_{1}(x)\sim\mathcal{N}(\mu_{1},\sigma^{2}\mathbf{I})$, $Q_{0}(x)\sim\mathcal{N}(\zeta_{0},\tau^{2}\mathbf{I})$, and $Q_{1}(x)\sim\mathcal{N}(\zeta_{1},\tau^{2}\mathbf{I})$, the Chernoff Difference is given as the absolute difference of Bhattacharyya distances:
$\text{CD}=\left|\frac{\|\mu_{0}-\mu_{1}\|_{2}^{2}}{8\sigma^{2}}-\frac{\|\zeta_{0}-\zeta_{1}\|_{2}^{2}}{8\tau^{2}}\right|.$ (11)
The derivation of this closed-form expression is provided in Appendix C and discussed in Nielsen (2022). This closed-form expression allows us to directly compute the CD from the definition of the distributions. Next, to incorporate our notion of privacy, we obtain the noisy variant of CD by leveraging the normal-sum theorem (Lemons and Langevin, 2002).
Lemma 4.
Under the same distributions as Lemma 3, for $\eta^{2}\geq 0$, the Noisy Chernoff Difference is:
$\widetilde{\text{CD}}_{\eta^{2}}=\left|\frac{\|\mu_{0}-\mu_{1}\|_{2}^{2}}{8(\sigma^{2}+\eta^{2})}-\frac{\|\zeta_{0}-\zeta_{1}\|_{2}^{2}}{8(\tau^{2}+\eta^{2})}\right|.$ (12)
To analyze this relationship in more depth, we examine the behavior of $\widetilde{\text{CD}}_{\eta^{2}}$ as we vary the noise parameter $\eta^{2}$. This leads to a central result of this work.
Theorem 1.
Suppose $P_{0}(x)\sim\mathcal{N}(\mu_{0},\sigma^{2}\mathbf{I})$, $P_{1}(x)\sim\mathcal{N}(\mu_{1},\sigma^{2}\mathbf{I})$, $Q_{0}(x)\sim\mathcal{N}(\zeta_{0},\tau^{2}\mathbf{I})$, and $Q_{1}(x)\sim\mathcal{N}(\zeta_{1},\tau^{2}\mathbf{I})$. Without loss of generality, we assume that $\|\mu_{0}-\mu_{1}\|_{2}\geq\|\zeta_{0}-\zeta_{1}\|_{2}$. There are three behaviors of the Noisy Chernoff Difference ($\widetilde{\text{CD}}_{\eta^{2}}$): (i) $\widetilde{\text{CD}}_{\eta^{2}}$ has a maximum point, (ii) $\widetilde{\text{CD}}_{\eta^{2}}$ has a maximum point and a reflection point (where $\widetilde{\text{CD}}_{\eta^{2}}=0$), (iii) $\widetilde{\text{CD}}_{\eta^{2}}$ is non-increasing. [^1] When $\|\mu_{0}-\mu_{1}\|_{2}=\|\zeta_{0}-\zeta_{1}\|_{2}$, $\widetilde{\text{CD}}_{\eta^{2}}$ will always fall into this case. The respective conditions for these three cases are given as follows:
  1. (i)
    $\frac{\|\zeta_{0}-\zeta_{1}\|^{2}_{2}}{\|\mu_{0}-\mu_{1}\|^{2}_{2}}<\frac{\tau^{2}}{\sigma^{2}}<\frac{\|\zeta_{0}-\zeta_{1}\|_{2}}{\|\mu_{0}-\mu_{1}\|_{2}}<1$, (Case 1: Maximum Point)
  2. (ii)
    $\frac{\tau^{2}}{\sigma^{2}}<\frac{\|\zeta_{0}-\zeta_{1}\|^{2}_{2}}{\|\mu_{0}-\mu_{1}\|^{2}_{2}}<1$, (Case 2: Maximum and Reflection)
  3. (iii)
    Neither condition (i) or (ii) hold. (Case 3: Non-increasing)
We provide a proof for this theorem in Appendix C. These three cases correspond to different scenarios where we have distinct relationships between fairness, privacy, and accuracy.
(a) Case 1: Chernoff Difference
(b) Case 1: Fairness Accuracy
(c) Case 2: Chernoff Difference
(d) Case 2: Fairness Accuracy
(e) Case 3: Chernoff Difference
(f) Case 3: Fairness Accuracy
Figure 2: (Case 1) (a) $\widetilde{\text{CD}}_{\eta^{2}}$ increases until the maximum point, which represents the worst fairness-accuracy trade-off we encounter by adding noise. (b) The slopes of the fairness-accuracy plots increase as we add noise until the $\widetilde{\text{CD}}_{\eta^{2}}$ reaches the maximum. Parameters used here are: $\mu_{0}=0,\mu_{1}=16.5,\sigma=2.43$ and $\zeta_{0}=0.5,\zeta_{1}=3.8,\tau=0.55$. (Case 2) (c) $\widetilde{\text{CD}}_{\eta^{2}}$ decays until it reaches 0. It then reflects back and follows a similar pattern to Case 1. The maximum occurs at a large $\eta^{2}$ (6.86) value and thus is not plotted (Appendix D.2). (d) The steepness of the fairness-accuracy plots decreases as we add noise and CD decreases. The intersecting lines show that in some cases, we can achieve better fairness for the same accuracy when we add noise. Parameters used here are: $\mu_{0}=-4.2,\mu_{1}=1.3,\sigma=3$ and $\zeta_{0}=0.3,\zeta_{1}=2.7,\tau=0.25$. (Case 3) (e) $\widetilde{\text{CD}}_{\eta^{2}}$ decays steadily over the entire positive $\eta^{2}$ regime. (f) Fairness-accuracy curves become flatter as $\widetilde{\text{CD}}_{\eta^{2}}$ decreases with the increasing noise. Parameters used here are: $\mu_{0}=-4.2,\mu_{1}=1.3,\sigma=0.85$ and $\zeta_{0}=0.6,\zeta_{1}=1.6,\tau=0.6$.
Experiments.
We now examine each case in Theorem 1, focusing on how the conditions relate to properties of the distributions and how different patterns in Chernoff Difference shape the interplay between fairness, privacy, and accuracy. Using synthetic Gaussian data, we demonstrate that Chernoff Difference effectively predicts the shape of fairness–accuracy curves.
For the synthetic Gaussian experiments, we use 1-D Gaussian distributions for $P_{0}(x),P_{1}(x),Q_{0}(x),$ and $Q_{1}(x)$ and sample 100,000 samples from each distribution to create a balanced dataset. We train a different Gaussian Naïve Bayes classifier for each of the groups (this provides the Bayes optimal classifier). We measure the overall accuracy of our classifier and quantify fairness by using the true positive rate disparity between groups (as this corresponds to the dominating exponent). To achieve fairness-accuracy trade-offs, we choose different prior beliefs of our labels for the Gaussian Naïve Bayes classifier. By perturbing the prior probabilities for the unprivileged group, we create fairness-accuracy curves for each of our settings [^2] A proxy for adjusting the threshold of the Bayes optimal classifier (Dutta et al., 2020)..
Case 1: Privacy Can Hurt Fairness.
In Case 1, $\widetilde{\text{CD}}_{\eta^{2}}$ grows as we increase the variance of our Gaussian noise until it reaches a maximum. After this point, additional noise leads to a decay in $\widetilde{\text{CD}}_{\eta^{2}}$ (see Figure 2(a)). In this case, $Q$ has means that are closer than those of $P$. At the same time, the variance of group $Q$’s distribution must fall within a specific range: it must be large enough (compared to $P$) so that $Q$ would be the unprivileged group (i.e., less separable), but also small enough that small amounts of added noise degrade $Q$’s separability more than $P$’s. In this regime, because $Q$ is already less separable, the gap in Chernoff Information between the groups increases as small noise is added. Once the noise becomes sufficiently large, it begins to affect $P$’s separability slightly more, marking a transition where the Chernoff Difference starts to decrease. This is validated through the respective fairness accuracy curves (Figure 2(b)). We observe that as we approach our maximum value of the Chernoff Difference, the slopes of our fairness-accuracy plots become steeper, highlighting a scenario in which adding privacy hurts both accuracy and fairness. We provide log fairness-accuracy plots in Appendix D to help discern how the shapes of the curves change.
Case 2 and Case 3: Privacy Can Help Fairness.
In Case 2, the Chernoff Difference decays to 0 before it starts to increase (Figure 2(c)). This occurs in a similar setting to Case 1, but where the variance of group $Q$ is much smaller. As a result, $Q$ becomes the privileged group, with better separability (i.e., higher Chernoff Information). As we add noise, the separability of $P$ and $Q$ both decrease, but $Q$ much faster than $P$, due to its smaller mean separation and lower variance. Because the privileged group loses separability faster, the Chernoff Difference decreases, effectively reducing unfairness, creating room for “free fairness.” This continues until the reflection point is reached where $\widetilde{\text{CD}}_{\eta^{2}}=0$. Beyond this point, the behavior mirrors that of Case 1. This reduction in unfairness is reflected in the fairness accuracy curves where curves rapidly flattens (Figure 2(d)). In fact, the curves flatten so quickly that the noisy curves begin to overlap with the clean fairness-accuracy curves. This shows that by adding privacy, we can achieve better fairness for the same accuracy—privacy gives free fairness! This emphasizes that, depending on the data distributions, there exist cases fairness and privacy are not at odds with each other, and in fact, they can work in tandem.
In Case 3, when neither condition (i) or (ii) from Theorem 1 hold, $\widetilde{\text{CD}}_{\eta^{2}}$ smoothly decays as we increase $\eta^{2}$ (Figure 2(e)). Here, group $Q$ exhibits a larger variance in addition to their smaller mean separation (compared to group $P$), leading to significantly less separability. As noise is added, the separability of $P$ is reduced at a rate faster than $Q$, however, the separability of $Q$ is initially too small for $P$ to catch up in the positive $\eta^{2}$ regime, leading to the non-increasing behavior. In Figure 2(f), we observe the fairness-accuracy curve flattening as the Chernoff Difference decays—similar to the trend seen in Case 2. However, in this scenario, we do not see any overlap between the fairness-accuracy curves, as the decay of the Chernoff Difference is slow. In this case, privacy may help the fairness-accuracy tradeoff (flatten the curve), but at a steep cost of accuracy.
5 Chernoff Information Neural Estimation (CINE)
Beyond isotropic Gaussian distributions, closed form solutions for Chernoff Information (CI) are only available for single-parameter members of the exponential family (Nielsen, 2013). On real-world data, computing CI requires density estimation, a long-standing research challenge in ML and statistics (Bishop and Nasrabadi, 2006). Here, we introduce Chernoff Information Neural Estimation (CINE), a novel algorithm that estimates CI from arbitrary distributions by recasting it as a density ratio estimation problem. To the best of our knowledge, this is the first method for estimating CI directly from real-world datasets. In order to develop our CI estimation algorithm, we first make the following crucial observation:
Observation 1.
Chernoff Information can be written as an optimization of an expectation
$C(P_{0},P_{1})=-\inf_{u\in(0,1)}\log\mathbb{E}_{x\sim P_{0}}\left[\left(\frac{P_{1}(x)}{P_{0}(x)}\right)^{u}\right].$ (13)
This observation yields a density ratio within the expectation. Accurately estimating the density ratio $\frac{P_{1}(x)}{P_{0}(x)}$ is easier than directly estimating $P_{1}(x)$ and $P_{0}(x)$ (Sugiyama et al., 2012; Vapnik, 2006), as shown theoretically in Kanamori et al. (2010). We leverage this insight to develop our algorithm.
Algorithm 2 Chernoff Information Neural Estimation via DRE-$\infty$ (Choi et al., 2022)
1:Distributions $\bm{P_{0}},\bm{P_{1}}$
[…]
6:return CICI
Background on neural density ratio estimation.
To estimate the density ratio, we adopt a state-of-the-art method known as the telescoping approach (Choi et al., 2022; Rhodes et al., 2020), which expresses the ratio $r(x)=\frac{P_{1}(x)}{P_{0}(x)}$ as a product of ratios between intermediate bridge distributions $p_{\lambda}(x)$. These bridges, indexed by $\lambda$, are created via an interpolation scheme between the original distributions (e.g., $p_{\lambda}(x)=\lambda P_{0}(x)+\sqrt{1-\lambda^{2}}P_{1}(x)$). This approach improves estimation accuracy by decomposing the overall ratio into smaller steps between distributions that are closer to each other and thus easier to estimate. Choi et al. (2022) consider the case where the number of bridges approaches infinity, allowing the log-density ratio to be expressed as an integral: $\log r(x)=\int_{1}^{0}\frac{\partial}{\partial\lambda}\log p_{\lambda}(x)\partial\lambda$. Observing the similarity between this telescoping process and diffusion models (Nichol and Dhariwal, 2021; Song et al., 2021), they propose training a neural network to recover the time score function $\frac{\partial}{\partial\lambda}p_{\lambda}(x)$, while the integral is computed using standard numerical methods.
We refer the reader to the original work (Choi et al., 2022) for full details of the estimator and Appendix F.1 for details regarding our implementation.
Once the neural density ratio estimator is trained, we compute the expectation $\mathbb{E}_{x\sim P_{0}}[(\frac{P_{1}(x)}{P_{0}(x)})^{u}]$ using a Monte Carlo method, and denote it as a function $g(u)$. Finally, the last step of estimating Chernoff Information is solving $\inf_{u\in(0,1)}g(u)$ in (13). We highlight that this optimization can be performed easily and efficiently due to the convex nature of $g$. Convexity follows from a result by Nielsen (2022):
Lemma 5 (Section 2.1 (Nielsen, 2022)).
The skewed Bhattacharyya distance, $\log(\int P_{0}(x)^{1-u}P_{1}(x)^{u}dx)$ is convex with respect to $u$.
We provide a full overview of CINE in Algorithm 2. The key insight of CINE is that by reformulating the definition of CI, we can leverage both state-of-the-art methods for density ratio estimation (Choi et al., 2022) and recent theoretical advances on CI (Nielsen, 2022). Finally, we prove the consistency of the proposed CINE algorithm (Appendix E):
Theorem 2.
Let $P_{0}$ and $P_{1}$ be distributions with common support and bounded density ratio. Then the Chernoff Information estimator, constructed using a consistent density ratio estimator, is itself consistent.
Figure 3: (Algorithm 2 performance on Gaussian Data) (a) In 2D, using distributions $\mathcal{N}(\textbf{0},\sigma^{2}\mathbf{I})$ and $\mathcal{N}(\textbf{1},\sigma^{2}\mathbf{I})$ Chernoff Information estimates remain accurate for different values of $\sigma^{2}$. (b) In 5D, using distributions $\mathcal{N}(\textbf{0},\frac{1}{2}\mathbf{I})$ and $\mathcal{N}(\textbf{1},\textbf{I})$ Chernoff Information estimation remains accurate, even as noise $(\eta^{2})$ is added. (c) Using distributions $\mathcal{N}(\textbf{0},\frac{1}{2}\mathbf{I})$ and $\mathcal{N}(\textbf{1},\textbf{I})$, estimations are accurate in lower dimensions before degrading.
5.1 Empirical Evaluation of CINE
Next, we validate CINE by comparing its results with both the closed-form solutions of CI and the algorithm proposed by Nielsen (2022). This evaluation demonstrates the correctness of our method and its scalability with input dimension. In all experiments, we used 10,000 samples per class to train CINE. More details about the setting can be found in Appendix F.2.
We first test our algorithm on Gaussian distributions that have a closed form solution for Chernoff Information (see Lemma 3). We construct 2D Gaussian distributions with the equal variance: $\mathcal{N}(\textbf{0},\sigma^{2}\mathbf{I})$ and $\mathcal{N}(\textbf{1},\sigma^{2}\mathbf{I})$. As observed in Figure 3(a), our estimator closely matches the closed form solution of across a wide range of variances ($\sigma^{2}$).
Next, to evaluate on distributions with no closed form solution, we compare our estimator to the multidimensional Gaussian estimator from Nielsen (2022). This estimator relies on the fact that the CI between high dimensional Gaussians, can be written as a skewed KL-divergence where the optimal skewing parameter can be found with a binary search algorithm [^3] This algorithm relies on knowledge that the distributions are Gaussian.. For experiments, we construct 5D Gaussian distributions $N(\mathbf{0},(\frac{1}{2}+\eta^{2})\mathbf{I})$ and $\mathcal{N}(\mathbf{1},\eta^{2}\mathbf{I})$. We then sweep across various noise values $\eta^{2}$ and show that our algorithm closely matches Nielsen’s estimator.
Finally, in Figure 3(c), we examine how the performance of CINE scales with input dimension. We construct Gaussian distributions, $N(\mathbf{0},\frac{1}{2}\mathbf{I})$ and $\mathcal{N}(\mathbf{1},\mathbf{I})$, with dimensions ranging from 5 to 40. Here we can see that with 10,000 samples per class, our estimator closely matches Nielsen’s up to 20 dimensions. Beyond 20 dimensions, CINE begins to deviate, slightly overestimating CI compared to Nielsen’s results and exhibiting higher variance. This higher dimensional degradation is a common problem when estimating information-theoretic measures from samples (Belghazi et al., 2018), and we further discuss these limitations in Section 7.
6 Fairness-Privacy Dynamic via CINE
Mixture of Gaussians.
To move beyond the isotropic Gaussian experiments we conduct experiments using a mixture of Gaussians. For each conditional distribution, we utilize two 2D Gaussian distributions to create the mixture. We consider 2 mixtures, Mixture 1 and Mixture 2 and repeat the experiments from Section 4 (for more details on these mixtures and experiments, please refer to Appendix F). Here we observe that we are able to recover cases very similar to the isotropic Gaussian experiments. Using Mixture 1 we demonstrate that the CD spikes as we increase the variance of the noise (Figure 5(a)), similar to the trend we observed in Case 1 with isotropic Gaussians. Additionally, we observe a steepening in the fairness accuracy curve similar to the previous case (Figure 5(b)). Using Mixture 2 we demonstrate that CD can quickly decay towards 0 like in Case 2 from the earlier setting (Figure 5(c)). Further we see that the fairness accuracy curves flatten as noise increases and we can even obtain free fairness (Figure 5(d)).
(a) Mixture 1: CD
(b) Mixture 1: Fairness-Acc
(c) Mixture 2: CD
(d) Mixture 2: Fairness-Acc
(e) Adult: CD
(f) Case 3: Fairness-Acc
(g) MNIST: CD
(h) MNSIT: Fairness-Acc
Figure 5: (Mixture 1) (a) $\widetilde{\text{CD}}_{\eta^{2}}$ spikes as $\eta^{2}$ increases. (b) Fairness-Accuracy curve becomes more steep resembling fairness-accuracy curves from Case 1 from isotropic Gaussians. (Mixture 2) (c) $\widetilde{\text{CD}}_{\eta^{2}}$ decreasing to 0 (d) Fairness-Accuracy curves flattening resembling fairness-accuracy curves from Case 2 from isotropic Gaussians. (Adult) (e) $\widetilde{\text{CD}}_{\eta^{2}}$ Remains Flat for Adult dataset. (f) Fairness-Accuracy slopes remain stable for Adult dataset. (MNIST) (g) $\widetilde{\text{CD}}_{\eta^{2}}$ decreases for experiments on MNIST representations. (h) Fairness-Accuracy Curves Flatten for MNIST representations.
Adult.
We next examine the Adult dataset (Becker and Kohavi, 1996) to investigate how Noisy CD can provide insights into the triad dynamic. For our task, we consider the standard income classification task and use sex as our protected attribute. For our features, we select all the continuous features from the dataset, giving us 6 dimensions. Finally, we scale all the data to have a 0 mean and unit variance. For full details on these experiments please refer to Appendix F.3.
First, as the variance of the additive Gaussian noise increases, the CD remains relatively flat (Figure 5(e)). We observe that there may be slight trends in CD, however, they appear to be too small to make meaningful differences in the fairness, privacy dynamic. This is reflected by the fairness accuracy curves (Figure 5(f)). We train split logistic regression classifiers and perturb the class weights for the disadvantaged group to create our fairness-accuracy curves. For more details, please refer to Appendix F.3. As observed in Section 4, the CD acts as a proxy for the slope of the fairness accuracy curves. Here, as the input noise varies, the slopes of the fairness accuracy curves remain relatively stable. Further, as the CD is consistently low, we observe very flat fairness accuracy curves. This indicates that privacy does not have a strong effect on the relationship between fairness and accuracy for these features.
MNIST.
To extend our analysis beyond continuous tabular data, we propose using CI to analyze image representations. To do this, we design a pseudo-fairness-accuracy task on lower-dimensional MNIST representations. To create the representations, we perform PCA on the MNIST data and extract 2 principal components. We then assign specific digits to represent each subgroup in our fairness-privacy-accuracy experiment. For $P_{0}$, we select the digit 0, $P_{1}$, we select the digit $2$, $Q_{0}$, we select the digit $3$ and $Q_{1}$ we select the digit 7. Using these subgroups, we repeat our experiments from the previous sections, measuring Noisy CD and generating pseudo-fairness-accuracy curves. For CD, we observe a steady decay as we increase the input perturbation noise (Figure 5(g)). This trend is reflected on the pseudo-fairness-accuracy curves in which we observe a flattening effect (Figure 5(h)). Thus, CD is able to capture the slope of the fairness-accuracy even on representations of more complex data.
7 Final Remark
This work focuses on input perturbation as the privacy mechanism, since it maps directly to Chernoff Information. However, this notion of privacy limits our analysis to continuous data. Extending our analysis to other private learning algorithms such as DP-SGD (Abadi et al., 2016) would be interesting future work. Our Chernoff information estimator marks a step toward bridging information theory and practical implementation, and we show that it performs well empirically. However, we leave a full theoretical analysis for future work. In this paper, we focus on the practically relevant regime of small to moderate dimensionality, in which we can obtain density ratio estimates with a reasonable number of samples. Exploring improvements in higher dimensions is an interesting future line of work. We additionally highlight applications to higher-dimensional settings via representations as a valuable future direction. Finally, the insights from our work may inform the design of future fair and private learning algorithms that account for the structure of the data distribution.
Acknowledgments
This work was supported by the National Science Foundation (NSF) under grant number 2341055.
Disclaimer.
This paper was prepared for informational purposes by the Global Technology Applied Research center of JPMorgan Chase & Co. This paper is not a product of the Research Department of JPMorgan Chase & Co. or its affiliates. Neither JPMorgan Chase & Co. nor any of its affiliates makes any explicit or implied representation or warranty and none of them accept any liability in connection with this paper, including, without limitation, with respect to the completeness, accuracy, or reliability of the information contained herein and the potential legal, compliance, tax, or accounting effects thereof. This document is not intended as investment research or investment advice, or as a recommendation, offer, or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction.
Appendix
Appendix A provides background on privacy, the input perturbation notion, and how noise connects to differential privacy. Appendix B provides more background on Chernoff Information and its connection to fairness. Appendix C reiterates useful Theorems and Lemmas from the Isotropic Gaussian examples as well as presents their proofs. Appendix D provides a comparison between the fairness-accuracy curves and the log-fairness-accuracy curves for the initial Gaussian experiments. Additionally it holds a more comprehensive overview of Case 2 in the Gaussian setting. Appendix E contains the proof of consistency for the Chernoff Information Neural Estimator. Appendix F holds experimental details. Appendix G contains additional experiments on the HSLS dataset. Appendix H provides a comparison of the fairness-accuracy curves and the log-fairness-accuracy curves for real world data. Finally, Appendix I provides a brief ablation over hyperparameter choices.
Appendix A Background on Privacy
In this work, we utilize the common notion of differential privacy.
Definition A.1 (Differential Privacy).
A randomized learning algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-differentially private if, for every pair of neighboring datasets $D$ and $D^{\prime}$ that differ in at most one element, and for every measurable subset $S$ of the output space of $\mathcal{A}$,
$\Pr\!\bigl[\mathcal{A}(D)\in S\bigr]\;\leq\;e^{\varepsilon}\,\Pr\!\bigl[\mathcal{A}(D^{\prime})\in S\bigr]+\delta.$ (14)
Throughout this work, we utilize the input perturbation notion of privacy (Algorithm 1). For input perturbation, the naive bound on privacy stems from the feature level notion of neighboring datasets.
Definition A.2 (Neighboring Datasets – Feature Level).
Two datasets $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ and $D^{\prime}=\{(x^{\prime}_{i},y_{i})\}_{i=1}^{n}$ are neighboring at the feature-level if they differ in the features of a single element only. That is,
$\exists\,j\in[n]\text{ such that }x_{j}\neq x^{\prime}_{j}\text{ but }y_{j}=y^{\prime}_{j},$
and for all $i\neq j$, $x_{i}=x^{\prime}_{i}$ and $y_{i}=y^{\prime}_{i}$.
The input perturbation notion of privacy leverages the Gaussian Mechanism to achieve differential privacy.
Definition A.3 (Gaussian Mechanism (Dwork et al., 2014)).
Given a function $f:\mathcal{D}\to\mathbb{R}^{d}$ with $\ell_{2}$-sensitivity $\Delta_{2}(f)=\sup_{D,D^{\prime}}\|f(D)-f(D^{\prime})\|_{2}$ over all neighboring datasets $D,D^{\prime}$, the Gaussian Mechanism with parameter $\eta>0$ outputs
$\mathcal{M}(D)=f(D)+\mathcal{N}(0,\eta^{2}\bm{I}_{d}),$
where $\mathcal{N}(0,\sigma^{2}\bm{I}_{d})$ is the $d$-dimensional Gaussian distribution with mean 0 and covariance $\eta^{2}\bm{I}_{d}$.
It follows that the Gaussian Mechanism can provide differential privacy
Theorem A.1 (Theorem A.1 (Dwork et al., 2014)).
Let $\varepsilon\in(0,1)$ be arbitrary. For $c^{2}\geq 2\ln(1.25/\delta)$ the Gaussian Mechanism with parameter $\eta\geq c\Delta_{2}(f)/\varepsilon$ is $(\varepsilon,\delta)$-differentially private.
The input perturbation notion of privacy is equivalent to applying the Gaussian Mechanism to the identity function on our dataset and allows us to derive a relationship between privacy parameters $(\varepsilon,\delta)$ and the noise parameter $\eta^{2}$.
Lemma A.1 (Dwork et al. (2014)).
Assume $\|\mathbf{x}\|_{2}\leq 1$ for all $\mathbf{x}\in\mathbb{R}^{d}$. When $\eta^{2}\geq\frac{8\log(1.25/\delta)}{\varepsilon^{2}}$, Algorithm 1 is $(\varepsilon,\delta$)-DP with respect to feature-level neighboring datasets (Definition A.2).
Proof.
Since $\|\mathbf{x}\|_{2}\leq 1$, the $\ell_{2}$-sensitivity of the identity function is bounded by $2$. Thus, by Theorem A.1, we get $(\varepsilon,\delta)$-differential privacy with noise $\eta^{2}\geq\frac{8\log(1.25/\delta)}{\varepsilon^{2}}$. This $(\varepsilon,\delta)$-differential privacy extends to Algorithm 1 by the postprocessing property of differential privacy. ∎
Recent works have explored tightening this bound and generalizing to the more common notion of neighboring datasets.
Definition A.4 (Neighboring Datasets).
Two datasets $D=\{(x_{i},y_{i})\}_{i=1}^{n}$ and $D^{\prime}=\{(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{n}$ are neighboring if they differ in the data of a single element. That is,
$\exists\,j\in[n]\text{ such that }(x_{j},y_{j})\neq(x^{\prime}_{j},y^{\prime}_{j}),$
and for all $i\neq j$, $(x_{i},y_{i})=(x^{\prime}_{i},y^{\prime}_{i})$.
For $T$ training steps and $n$ samples, Kang et al. (2020) show that the noise bound becomes $O(\frac{G^{2}T\log}{n(n-1)\sqrt{\Delta}\varepsilon^{2}})$ for a $G$-Lipschitz and $\Delta$-strongly convex loss function . Fukuchi et al. (2017), show that for a quadratic loss function the noise bound becomes $O(\frac{G^{2}\log(16/\delta))+\varepsilon}{n\varepsilon^{2}})$. In all cases, $\eta^{2}\propto\frac{\log(1/\delta)}{\varepsilon^{2}}$. Thus, throughout this paper we utilize $\eta^{2}$ as our privacy parameter, since it provides a simple and standard expression directly linked to $(\varepsilon,\delta)$ via the Gaussian mechanism, allowing us to avoid restricting our analysis with assumptions on the learning algorithm. While tighter bounds can be derived, they do not affect the qualitative behavior of our results.
Appendix B Background on Chernoff Information
B.1 Bounding the Error of Bayes Optimal Classifier
Lemma B.1 ((Nielsen, 2011)).
For hypotheses $P_{0}(x)$ under $Y=0$ and $P_{1}(x)$ under $Y=1$, Chernoff Information bounds the error of the Bayes Optimal Classifier $H$:
$\Pr[H(X)\neq Y]\leq e^{-C(P_{0},P_{1})}.$
Proof.
Note: Most of this proof is from Nielsen (2011). We have included it for completeness.
Consider the binary classification setting. Let $\Pr[Y=0]=w_{0}>0$ and let $\Pr[Y=1]=w_{1}=1-w_{0}>0$. The class probabilities can be defined as $p_{0}(x)=\Pr[x|Y=0]$ and $p_{1}(x)=\Pr[x|Y=1]$. Now, consider the Bayes decision rule in which $H(x)$ classifies as $\hat{Y}=0$ if $Pr[Y=0|x]>Pr[Y=1|x]$ and $\hat{Y}=1$ otherwise. Using Bayes rule, $\Pr[Y=i|x]=\frac{\Pr[Y=i]\Pr[x|Y=i]}{\Pr[x]}=\frac{w_{i}p_{i}(x)}{p(x)}$ for $i\in\{0,1\}$. Now, we can obtain the probability of error:
[…]
Thus, the Bayes decision rule minimizes by principle the average probability of error:
$E^{*}=\int\Pr[\text{Error}|x]p(x)dx=\int\min\{\Pr[Y=0|x],\Pr[Y=1|x]\}p(x)dx.$
Now, to upper bound this Bayes error, we can utilize knowledge that, for $a,b>0$, $\min\{a,b\}\leq a^{\alpha}b^{1-\alpha}~~\forall\alpha\in(0,1)$. This implies that:
$E^{*}\leq w_{0}^{\alpha}w_{1}^{1-\alpha}\int p_{0}(x)^{\alpha}p_{1}(x)^{1-\alpha}dx.$
Now, let $c_{\alpha}(p_{0}:p_{1})=\int p_{0}(x)^{\alpha}p_{1}(x)^{1-\alpha}dx$. This term is referred to as the Chernoff $\alpha$-coefficient. Since the above inequality holds for all $\alpha\in(0,1)$, the best exponent for upper bounding the Bayes error corresponds to the optimal Chernoff $\alpha$-coefficient:
$c^{*}(p_{0}:p_{1})=c_{\alpha^{*}}(p_{0}:p_{1})=\inf_{\alpha\in(0,1)}\int p_{0}(x)^{\alpha}p_{1}(x)^{1-\alpha}dx$
The Chernoff Coefficient gives a measure of similarity which yields CI:
$C(p_{0},p_{1})=-\log\inf_{\alpha\in(0,1)}\int p_{0}(x)^{\alpha}p_{1}(x)^{1-\alpha}dx.$
CI yields the best achievable exponent for a Bayesian probability of error:
$E^{*}\leq w_{0}^{\alpha}w_{1}^{1-\alpha}e^{-C(p_{0},p_{1})}.$
Lemma 1 follows as $w_{0},w_{1}<1$. ∎
B.1.1 Tightness of Bound
CI provides an asymptotically tight bound on the error exponent, i.e.,
$\lim_{n\rightarrow\infty}-\frac{1}{n}\log P_{e}^{(n)}=C(P_{0},P_{1}),$
where $n$ is the number of samples (Cover, 1999). For instance, for Gaussian distributions, we can explicitly write:
$c_{1}e^{-nC(P_{0},P_{1})}\leq P_{e}^{(n)}\leq c_{2}e^{-nC(P_{0},P_{1})},$
which shows that CI provides an exponent-tight bound ($c_{1}$ and $c_{2}$ are constants with a closed-form expression).
B.2 Connection to Fairness Metrics
Next, we provide an overview of Chernoff Information and its connection to fairness following results from Dutta et al. (2020).
Consider the binary classification task ($Y\in\{0,1\}$) with a binary protected attribute ($S\in\{0,1\}$). Consider the standard Bayes optimal classification rule of split classifiers: $H_{s}(x)=\frac{\Pr[Y=1|S=s,x]}{\Pr[Y=0|S=s,x]}\geq\tau_{s}$ in which the classifier predicts positive if the ratio of the conditionals is greater than a threshold $\tau_{s}$.
Lemma B.2 (Lemma 2 from Dutta et al. (2020)).
The Chernoff exponent of the probability of error of the Bayes optimal classifier is given by the Chernoff Information:
$E_{e}=C(P_{0},P_{1})=-\inf_{u\in(0,1)}\log(\int P_{0}(x)^{1-u}P_{1}(x)^{u}dx).$ (15)
As highlighted in Lemma 2, this Chernoff exponent bounds the probability of error of the Bayes optimal classifier. We now show a connection between false positive and false negative error rates and Chernoff Information.
Definition B.1 (Chernoff Exponents—Definition 1 from Dutta et al. (2020)).
The Chernoff exponents of the False positive error and False negative error are defined as
$E_{\text{FP}}(\tau_{s})=\sup_{u>0}(u\tau_{s}-\Lambda_{0}(u))$ (16)
$E_{\text{FN}}(\tau_{s})=\sup_{u<0}(u\tau_{s}-\Lambda_{1}(u))$ (17)
where $\Lambda_{0}$ and $\Lambda_{1}$ are the log-generating functions.
Similarly, these Chernoff exponents bound the probability of false positive and false negative errors.
Lemma B.3 (Chernoff Bound - Lemma 1 from Dutta et al. (2020)).
The Chernoff exponents of false positive error and false negative error bound the probability of false positive and false negative error
$P_{\text{FP}}\leq e^{-E_{FP}(\tau_{s})}$ (18)
$P_{\text{FN}}\leq e^{-E_{FN}(\tau_{s})}$ (19)
Under the equal prior condition, the overall Chernoff exponent can be written as simple function of the false positive and false negative exponents.
Definition B.2 (Definition 2 from Dutta et al. (2020)).
Suppose $\Pr[S=0]=\Pr[S=1]$ and $\Pr[Y=1|S=s]=\Pr[Y=0|S=s]$ for all $s$. The Chernoff exponent of the overall probability of error $P_{e}$ is defined as
$E_{e}=\min\{E_{FP}(\tau_{s}),E_{FN}(\tau_{s})\}$ (20)
Thus, Chernoff Difference naturally measures disparities in error exponents across groups (False positive or False negative exponents).
$\text{CD}=|\min\{E_{FP}(\tau_{P}),E_{FN}(\tau_{P})\}-\min\{E_{FP}(\tau_{Q}),E_{FN}(\tau_{Q})\}|$ (21)
In the asymptotic regime, CD expresses disparities in terms of the maximum error rate on a log-scale, capturing a geometric difference rather than the arithmetic difference used in common fairness metrics such as Equal Opportunity. Next, we demonstrate how Chernoff Difference relates to the standard arithmetic difference.
Theorem B.1.
Let $M_{p}=\max\{P_{FP}^{P},P_{FN}^{P}\}$ and $M_{q}=\max\{P_{FP}^{Q},P_{FN}^{Q}\}$. For all constants $0<c<C$ with $c\leq M_{P},M_{Q}\leq C$, we have
$c\left|\log M_{p}-\log M_{Q}\right|\leq\left|M_{p}-M_{Q}\right|\leq C\left|\log M_{p}-\log M_{Q}\right|.$ (22)
Proof.
Since $M_{p},M_{q}\in[c,C]$ with $c>0$, the Mean Value Theorem implies
$\log M_{p}-\log M_{q}=\log^{\prime}(\xi)(M_{p}-M_{q})=\frac{1}{\xi}(M_{p}-M_{q})$ (23)
for some $\xi$ between $M_{p}$ and $M_{q}$. Thus
$|M_{p}-M_{q}|=\xi\,|\log M_{p}-\log M_{q}|.$ (24)
Since $\xi\in[c,C]$, we have
$c\,|\log M_{p}-\log M_{q}|\;\leq\;|M_{p}-M_{q}|\;\leq\;C\,|\log M_{p}-\log M_{q}|.$ (25)
When we break the equal prior assumption the Chernoff Information continues to capture error exponents, however skewed by the class priors.
Definition B.3 (Section E.1 (Dutta et al., 2020)).
Let $\pi_{0}=P[Y=0]$ and $\pi_{1}=P[Y=1]$. The Chernoff exponent of the overall probability of error $P_{e}$ is defined as
$E_{e}=\min\{E_{FP}(\tau_{s})-\log 2\pi_{0},E_{FN}(\tau_{s})-\log 2\pi_{1}\}$ (26)
Thus, the Chernoff Difference continues to evaluate the difference of the error exponents across groups. In the case where where the exponents align as false negative exponents, we again achieve a similar value to equal opportunity (although skewed by the log-ratio of the priors).
[…]
Appendix C Isotropic Gaussian Proofs
Lemma C.1 (Restatement of Lemma 3).
When $P_{0}(x)\sim\mathcal{N}(\mu_{0},\sigma^{2}\mathbf{I})$, $P_{1}(x)\sim\mathcal{N}(\mu_{1},\sigma^{2}\mathbf{I})$, $Q_{0}(x)\sim\mathcal{N}(\zeta_{0},\tau^{2}\mathbf{I})$, and $Q_{1}(x)\sim\mathcal{N}(\zeta_{1},\tau^{2}\mathbf{I})$, the Chernoff Difference is given as:
$CD=\left|\frac{\|\mu_{0}-\mu_{1}\|_{2}^{2}}{8\sigma^{2}}-\frac{\|\zeta_{0}-\zeta_{1}\|_{2}^{2}}{8\tau^{2}}\right|.$
Proof.
Recall the definition of Chernoff Difference and Chernoff Information
$CD$ $=|C(P_{0},P_{1})-C(Q_{0},Q_{1})|$ (28)
$=|\min_{u\in(0,1)}\log\int P_{0}(x)^{u}P_{1}(x)^{1-u}dx-\min_{v\in(0,1)}\log\int Q_{0}(x)^{v}Q_{1}(x)^{1-v}dx|.$ (29)
Now, following a result from Dutta et al. (2020), we see that for $P_{0}\sim\mathcal{N}(\mu_{0},\sigma^{2}\mathbf{I})$, $P_{1}\sim\mathcal{N}(\mu_{1},\sigma^{2}\mathbf{I})$.
$\log\int P_{0}(x)^{u}P_{1}(x)^{1-u}dx$ (30)
$=\log\int e^{-\frac{u}{2\sigma^{2}}\left((x-\mu_{1})^{T}(x-\mu_{1})-(x-\mu_{0})^{T}(x-\mu_{0})\right)}P_{0}(x)\,dx$ (31)
$=\log e^{-\frac{u}{2\sigma^{2}}\left(\mu_{1}^{T}\mu_{1}-\mu_{0}^{T}\mu_{0}\right)}\int e^{-\frac{u}{2\sigma^{2}}\left(-2x^{T}(\mu_{1}-\mu_{0})\right)}P_{0}(x)\,dx$ (32)
$=\log e^{-\frac{u}{2\sigma^{2}}\left(\mu_{1}^{T}\mu_{1}-\mu_{0}^{T}\mu_{0}\right)}e^{-\frac{u}{2\sigma^{2}}\left(-2\mu_{0}^{T}(\mu_{1}-\mu_{0})\right)}e^{\frac{u^{2}}{2\sigma^{2}}\left(\|\mu_{1}-\mu_{0}\|_{2}^{2}\right)}$ (33)
$=\log e^{-\frac{u}{2\sigma^{2}}\left(\|\mu_{1}-\mu_{0}\|_{2}^{2}\right)}e^{\frac{u^{2}}{2\sigma^{2}}\left(\|\mu_{1}-\mu_{0}\|_{2}^{2}\right)}$ (34)
$=\frac{u(u-1)}{2\sigma^{2}}\|\mu_{1}-\mu_{0}\|_{2}^{2}.$ (35)
Now, we can compute the derivative to minimize with respect to $u$:
$\frac{d}{du}\frac{u(u-1)}{2\sigma^{2}}\|\mu_{1}-\mu_{0}\|_{2}^{2}=\frac{2u-1}{2\sigma^{2}}\|\mu_{1}-\mu_{0}\|_{2}^{2}.$ (36)
The derivative is 0, when $u=0.5$. This critical point is a minimum following a first derivative test. Thus,
$C(P_{0},P_{1})=\frac{\|\mu_{0}-\mu_{1}\|_{2}^{2}}{8\sigma^{2}}.$ (37)
When these distributions are 1-dimensional, they reduce to the closed form derived in Nielsen (2022). A similar approach can be done for $Q_{0}(x)$ and $Q_{1}(x)$. Thus, the CD for Gaussian distributions is defined as
$CD=|\frac{\|\mu_{0}-\mu_{1}\|_{2}^{2}}{8\sigma^{2}}-\frac{\|\zeta_{0}-\zeta_{1}\|_{2}^{2}}{8\tau^{2}}|.$ (38)
Theorem C.1 ( Restatement of Theorem 1).
Suppose $P_{0}(x)\sim\mathcal{N}(\mu_{0},\sigma^{2}\mathbf{I})$, $P_{1}(x)\sim\mathcal{N}(\mu_{1},\sigma^{2}\mathbf{I})$, $Q_{0}(x)\sim\mathcal{N}(\zeta_{0},\tau^{2}\mathbf{I})$, and $Q_{1}(x)\sim\mathcal{N}(\zeta_{1},\tau^{2}\mathbf{I})$. Without loss of generality, we assume that $\|\mu_{0}-\mu_{1}\|_{2}\geq\|\zeta_{0}-\zeta_{1}\|_{2}$. There are three behaviors of the Noisy Chernoff Difference ($\widetilde{\text{CD}}_{\eta^{2}}$): (i) $\widetilde{\text{CD}}_{\eta^{2}}$ has a maximum point, (ii) $\widetilde{\text{CD}}_{\eta^{2}}$ has a maximum point and a reflection point (where $\widetilde{\text{CD}}_{\eta^{2}}=0$), (iii) $\widetilde{\text{CD}}_{\eta^{2}}$ is non-increasing. [^4] When $\|\mu_{0}-\mu_{1}\|_{2}=\|\zeta_{0}-\zeta_{1}\|_{2}$, $\widetilde{\text{CD}}_{\eta^{2}}$ will always fall into this case. The respective conditions for these three cases are given as follows:
  1. (i)
    $\frac{\|\zeta_{0}-\zeta_{1}\|^{2}_{2}}{\|\mu_{0}-\mu_{1}\|^{2}_{2}}<\frac{\tau^{2}}{\sigma^{2}}<\frac{\|\zeta_{0}-\zeta_{1}\|_{2}}{\|\mu_{0}-\mu_{1}\|_{2}}<1$, (Case 1: Maximum Point)
  2. (ii)
    $\frac{\tau^{2}}{\sigma^{2}}<\frac{\|\zeta_{0}-\zeta_{1}\|^{2}_{2}}{\|\mu_{0}-\mu_{1}\|^{2}_{2}}<1$, (Case 2: Maximum and Reflection)
  3. (iii)
    Neither condition (i) or (ii) hold. (Case 3: Non-increasing)
Proof.
Recall the definition of noisy Chernoff Difference. We define a signed noisy Chernoff Difference $s\widetilde{\text{CD}}_{\eta^{2}}$ such that $|s\widetilde{\text{CD}}_{\eta^{2}}|=\widetilde{\text{CD}}_{\eta^{2}}$. Thus,
$s\widetilde{\text{CD}}_{\eta^{2}}=\frac{1}{8(\tau^{2}+\eta^{2})}\|\zeta_{0}-\zeta_{1}\|^{2}_{2}-\frac{1}{8(\sigma^{2}+\eta^{2})}\|\mu_{0}-\mu_{1}\|^{2}_{2}.$ (39)
Case 1 and 2.
Let $p=\|\mu_{0}-\mu_{1}\|_{2}$ and let $q=\|\zeta_{0}-\zeta_{1}\|_{2}$. First, we can analyze the positive $\eta^{2}$ regime for a critical point. To find potential critical points, consider the derivative of $s\widetilde{\text{CD}}_{\eta^{2}}$.
$\frac{\partial}{\partial\eta^{2}}s\widetilde{\text{CD}}_{\eta^{2}}=\frac{p^{2}}{8(\sigma^{2}+\eta^{2})^{2}}-\frac{q^{2}}{8(\tau^{2}+\eta^{2})^{2}}=\frac{p^{2}(\tau^{2}+\eta^{2})^{2}-q^{2}(\sigma^{2}+\eta^{2})^{2}}{8(\sigma^{2}+\eta^{2})^{2}(\tau^{2}+\eta^{2})^{2}}$ (40)
Now, the potential critical points will be $\eta^{2}$ where $s\widetilde{\text{CD}}_{\eta^{2}}^{\prime}=0$. That is $0=p^{2}(\tau^{2}+\eta^{2})^{2}-q^{2}(\sigma^{2}+\eta^{2})^{2}$. Thus, the critical points are:
$\eta_{0}^{2}=\frac{\sigma^{2}q-\tau^{2}p}{p-q}$ (41)
$\eta_{1}^{2}=\frac{-\sigma^{2}q-\tau^{2}p}{p+q}$ (42)
However, we observe that $\eta_{1}^{2}$ is always negative, thus the only useful critical point in the postive $\eta^{2}$ region is $\eta_{0}^{2}$. By our assumption, we know that $p-q\geq 0$. First, we observe that there is no critical point when $p=q$ as $\eta_{0}^{2}$ does not exist. Thus, $\eta_{0}^{2}$ is positive when the following conditions hold.
$\sigma^{2}q-\tau^{2}p>0$ (43)
$\frac{\tau^{2}}{\sigma^{2}}<\frac{\|\zeta_{0}-\zeta_{1}\|_{2}}{\|\mu_{0}-\mu_{1}\|_{2}}<1$ (44)
Next, we will show that this critical point is always a maximum in the positive $\eta^{2}$ regime.
First, consider the second derivative of the signed noisy Chernoff difference.
$\frac{\partial^{2}}{\partial(\eta^{2})^{2}}s\widetilde{\text{CD}}_{\eta^{2}}=\frac{q^{2}}{4(\tau^{2}+\eta^{2})^{3}}-\frac{p^{2}}{4(\sigma^{2}+\eta^{2})^{3}}.$ (45)
Plugging in the relevant critical point we observe that
$\frac{\partial^{2}}{\partial(\eta^{2})^{2}}s\widetilde{\text{CD}}_{\eta_{0}^{2}}$ $=\frac{q^{2}}{4(\tau^{2}+\eta_{0}^{2})^{3}}-\frac{p^{2}}{4(\sigma^{2}+\eta_{0}^{2})^{3}}=\frac{q^{2}}{4(\frac{q(\sigma^{2}-\tau^{2})}{p-q})^{3}}-\frac{p^{2}}{4(\frac{p(\sigma^{2}-\tau^{2})}{p-q})^{3}}$ (46)
$=\frac{q^{2}(p-q)^{3}}{4q^{3}(\sigma^{2}-\tau^{2})^{3}}-\frac{p^{2}(p-q)^{3}}{4p^{3}(\sigma^{2}-\tau^{2})^{3}}=\frac{(p-q)^{4}}{4p^{3}q^{3}(\sigma^{2}-\tau^{2})^{3}}.$ (47)
Now, we know that $\sigma^{2}>\tau^{2}$ so the second derivative of this critical point of $s\widetilde{\text{CD}}_{\eta^{2}}$ must be a minimum. However, the goal is to examine the behavior of $\widetilde{\text{CD}}_{\eta^{2}}$. So, we can show that this is critical point is a maximum of $\widetilde{\text{CD}}_{\eta^{2}}$, by showing that $s\widetilde{\text{CD}}_{\eta_{0}^{2}}$ is negative. By plugging in the critical point, we can observe that it is a maximum of $\widetilde{\text{CD}}_{\eta^{2}}$.
$s\widetilde{\text{CD}}_{\eta_{0}^{2}}$ $=\frac{q^{2}}{8(\tau^{2}+\eta_{0}^{2})}-\frac{p^{2}}{8(\sigma^{2}+\eta_{0}^{2})}=\frac{q^{2}}{8(\tau^{2}+\frac{\sigma^{2}q-\tau^{2}p}{p-q})}-\frac{p^{2}}{8(\sigma^{2}+\frac{\sigma^{2}q-\tau^{2}p}{p-q})}$ (48)
$=\frac{(p-q)}{8}(\frac{q}{\sigma^{2}-\tau^{2}}-\frac{p}{\sigma^{2}-\tau^{2}})=\frac{(p-q)(q-p)}{8(\sigma^{2}-\tau^{2})}$ (49)
Now, this value is always negative as we know $p>q$ and $\sigma^{2}>\tau^{2}$. Thus, the positive critical point is a maximum. Now, we can analyze the positive $\eta^{2}$ regime for a reflection point.
$s\widetilde{\text{CD}}_{\eta^{2}}=\frac{q^{2}}{8(\tau^{2}+\eta^{2})}-\frac{p^{2}}{8(\sigma^{2}+\eta^{2})}$ (50)
$8q^{2}(\sigma^{2}+\eta^{2})-8p^{2}(\tau^{2}+\eta^{2})=0$ (51)
$\eta^{2}=\frac{q^{2}\sigma^{2}-p^{2}\tau^{2}}{p^{2}-q^{2}}=\frac{\|\zeta_{0}-\zeta_{1}\|_{2}^{2}\sigma^{2}-\|\mu_{0}-\mu_{1}\|_{2}^{2}\tau^{2}}{\|\mu_{0}-\mu_{1}\|_{2}^{2}-\|\zeta_{0}-\zeta_{1}\|_{2}^{2}}.$ (52)
Now, the denominator of this is always positive, thus, $\eta^{2}$ is positive when the following holds:
$\|\zeta_{0}-\zeta_{1}\|_{2}^{2}\sigma^{2}-|\mu_{0}-\mu_{1}\|_{2}^{2}\tau^{2}>0$ (53)
$\frac{\tau^{2}}{\sigma^{2}}<\frac{\|\zeta_{0}-\zeta_{1}\|^{2}_{2}}{\|\mu_{0}-\mu_{1}\|^{2}_{2}}<1.$ (54)
Case 3.
Finally, we can examine the behavior when none of these conditions hold. Suppose $\frac{\tau^{2}}{\sigma^{2}}\geq\frac{\|\zeta_{0}-\zeta_{1}\|_{2}}{\|\mu_{0}-\mu_{1}\|_{2}}=\frac{q}{p}$. We can analyze the sign of the numerator of $s\widetilde{\text{CD}}_{\eta^{2}}^{\prime}$ by writing it as
$p^{2}(\tau^{2}+\eta^{2})^{2}-q^{2}(\sigma^{2}+\eta^{2})^{2}$ (55)
$=(p^{2}(\tau^{2}+\eta^{2})-q^{2}(\sigma^{2}+\eta^{2}))(p^{2}(\tau^{2}+\eta^{2})+q^{2}(\sigma^{2}+\eta^{2})).$ (56)
Thus, to analyze the sign, we can analyze $p^{2}(\tau^{2}+\eta^{2})-q^{2}(\sigma^{2}+\eta^{2})$. We observe
$p^{2}(\tau^{2}+\eta^{2})-q^{2}(\sigma^{2}+\eta^{2})=p^{2}\tau^{2}-q^{2}\sigma^{2}+\eta^{2}(p^{2}-q^{2}).$ (57)
Now, from our initial assumption, $p^{2}-q^{2}\geq 0$. From the assumption that $\frac{\tau^{2}}{\sigma^{2}}\geq\frac{q}{p}$, $p^{2}\tau^{2}-q^{2}\sigma^{2}\geq 0$. Thus, the sign is always positive. Now, to show that $\widetilde{\text{CD}}_{\eta^{2}}$ is non-increasing, we will show $s\widetilde{\text{CD}}_{\eta^{2}}$ is not positive.
$s\widetilde{\text{CD}}_{\eta^{2}}$ $=\frac{q^{2}}{8(\tau^{2}+\eta^{2})}-\frac{p^{2}}{8(\sigma^{2}+\eta^{2})}=\frac{q^{2}(\sigma^{2}+\eta^{2})-p^{2}(\tau^{2}+\eta^{2})}{8((\tau^{2}+\eta^{2}))(\sigma^{2}+\eta^{2})}$ (58)
Now, we analyze $q^{2}(\sigma^{2}+\eta^{2})-p^{2}(\tau^{2}+\eta^{2})$. We can see that this is equivalent to $q^{2}\sigma^{2}-p^{2}\tau^{2}+\eta^{2}(q^{2}-p^{2})$. From our initial assumption, $q^{2}-p^{2}\leq 0$. From the assumption that $\frac{\tau^{2}}{\sigma^{2}}\geq\frac{q}{p}$, $q^{2}\sigma^{2}-p^{2}\tau^{2}\leq 0$. Thus, $s\widetilde{\text{CD}}_{\eta^{2}}$ is always negative and $\widetilde{\text{CD}}_{\eta^{2}}$ is non-increasing over the positive $\eta^{2}$ regime. ∎
Appendix D Supplemental Gaussian Figures
D.1 Fairness-Accuracy vs Log Fairness-Accuracy
Figure D.1: (Case 1: Privacy Hurts Fairness) $\mu_{0}=0,\mu_{1}=16.5,\sigma=2.43$, $\zeta_{0}=0.5,\zeta_{1}=3.8,\tau=0.55$. (a) Fairness-Accuracy Curve. (b) Log Fairness-Accuracy Curve. We observe a steepening effect.
Figure D.2: (Case 2: Privacy Can Give Free Fairness) $\mu_{0}=-4.2,\mu_{1}=1.3,\sigma=3$, $\zeta_{0}=0.3,\zeta_{1}=2.7,\tau=0.25$. (a) Fairness-Accuracy Curve. (b) Log Fairness-Accuracy Curve. We observe a flattening effect.
Figure D.3: (Case 3: Triple Trade-off) $\mu_{0}=-4.2,\mu_{1}=1.3,\sigma=0.85$, $\zeta_{0}=0.6,\zeta_{1}=1.6,\tau=0.6$. (a) Fairness-Accuracy Curve. (b) Log Fairness-Accuracy Curve. We observe a much slower flattening effect.
D.2 Case 2: More detail
Figure D.4: (Case 2: Privacy Can Give Free Fairness) $\mu_{0}=-4.2,\mu_{1}=1.3,\sigma=3$, $\zeta_{0}=0.3,\zeta_{1}=2.7,\tau=0.25$. (a) Initial plot of $\widetilde{\text{CD}}_{\eta^{2}}$. (b) Full plot showing presence of maximum point.
Appendix E Consistency
E.1 Setting
Let $r(x)=\frac{P_{1}(x)}{P_{0}(x)}$ be the density ratio and let $f(u)=-\log\mathbb{E}_{x\sim P_{0}}[r(x)^{u}]$. We also define the corresponding empirical estimators $r_{n}(x)$ and $f_{n}(u)=-\log\frac{1}{n}\sum r_{n}(x_{i})^{u}$. We operate under the assumption that $P_{0}$ and $P_{1}$ share a support, $r(x)$ is bounded i.e., $r(x)\in[c,C]$ for some constants $0<c<C<\infty$.
We further assume that $r_{n}(x)$ is a consistent estimator of $r(x)$ and is bounded via a clipping mechanism i.e., $r_{n}(x)\in[c,C]$. Clipping is typically applied after estimation to prevent numerical instability and prevent estimates from growing too large or small. This clipping does not affect consistency under our assumptions if the clipping bound is wider than the true bounds on the density ratio. Finally, we assume that the estimator $r_{n}(x)$ is trained on its own samples, independent of the evaluation points.
E.2 Proofs
We will use Newey-McFadden’s main consistency theorem to show consistency of the Chernoff Information estimator.
Theorem E.1 (Newey-McFadden Main Consistency Theorem, Theorem 2.1 (Newey and McFadden, 1994)).
If there is a function f such that (i) $f(u)$ is uniquely maximized at $u^{*}$; (ii) $U$ is compact; (iii) $f$ is continuous; (iv) $f_{n}(u)$ converges uniformly in probability to $f(u)$, then $\hat{u}\rightarrow u^{*}$.
We begin with a series of lemmas before completing the proof of our theorem. First, we are able to show compactness via:
Lemma E.1.
We can write Chernoff Information as $\inf_{[0,1]}\log(\int P_{0}(x)^{1-u}P_{1}(x)^{u}dx)$.
Proof.
Computing the infimum over the open interval (0,1) and closed interval [0,1] yield the same result following the strict convexity of $f$. We point to Remark 2 from Nielsen (2022) for a discussion of the boundary conditions. In fact, in some references, Chernoff Information is presented using this closed interval (Cover, 1999). This provides a compactness property that can be utilized in the following lemmas. ∎
Lemma E.2.
$\{f_{n}\}$ converges pointwise in probability for $u\in[0,1]$
Proof.
First, we see that we can decompose the following difference into two terms,
$\frac{1}{n}\sum r_{n}(x_{i})^{u}-\mathbb{E}_{x\sim P_{0}}[r(x)^{u}]=(\frac{1}{n}\sum r_{n}(x_{i})^{u}-\mathbb{E}_{x\sim P_{0}}[r_{n}(x)^{u}])+(\mathbb{E}_{x\sim P_{0}}[r_{n}(x)^{u}]-\mathbb{E}_{x\sim P_{0}}[r(x)^{u}])$ (59)
Next, we can see that the first term $\frac{1}{n}\sum r_{n}(x_{i})^{u}-\mathbb{E}_{x\sim P_{0}}[r_{n}(x)^{u}]$ converges to 0 almost surely (a.s.) by the law of large numbers.
Since $r_{n}$ is consistent and uniformly bounded in $[c,C]$, there exists $\widetilde{r}_{n}(x)^{u}$ and $\widetilde{r}(x)^{u}$ defined on another space such that $\widetilde{r}_{n}(x)^{u}$ has the same distribution as $r_{n}(x)^{u}$, $\widetilde{r}(x)^{u}$ has the same distribution as $r(x)^{u}$, and $\widetilde{r}_{n}(x)^{u}$ converges to $\widetilde{r}(x)^{u}$ a.s. (via the Continuous Mapping Theorem and Skorohod’s Representation Theorem). The Dominated Convergence Theorem then implies $\mathbb{E}[\widetilde{r}_{n}(x)^{u}]-\mathbb{E}[\widetilde{r}(x)^{u}]$ converges to 0. Since the two sequences share the same distributions, convergence in the Skorokhod space implies convergence in the original space. It follows that $\mathbb{E}[r_{n}(x)^{u}]-\mathbb{E}[r(x)^{u}]$ converges to 0.
Since both terms converge to 0, their sum converges a.s. to 0 and hence in probability to 0. Thus, we have $\frac{1}{n}\sum r_{n}(x_{i})^{u}-\mathbb{E}[r(x)^{u}]\rightarrow 0$ in probability. By applying the continuous mapping theorem, we have $\log(\frac{1}{n}\sum r_{n}(x_{i})^{u})+\log(\mathbb{E}[r(x)^{u}])\rightarrow 0$. Thus, we have $\{f_{n}\}$ converges pointwise in probability. ∎
Lemma E.3.
$\{f_{n}\}$ is stochastically equicontinuous on $[0,1]$
Proof.
First, recall the density ratio $r(x)$ is bounded between $[c,C]$. Via clipping we assume the estimator $r_{n}(x)$ follows the same bound. Now, for any $x$, the function $u\rightarrow r_{n}(x)^{u}$ has a bounded derivative, $r_{n}(x)^{u}\log r_{n}(x)$, that only depends on constants $c$ and $C$, thus is Lipschitz in $u$. It follows that the average
$\frac{1}{n}\sum_{i=1}^{n}r_{n}(x_{i})^{u}$
is Lipschitz in $u$ again with the constant dependent only on $c$ and $C$. Now, since $\log(\cdot)$ is Lipschitz on the relevant range (away from 0), it follows from the compositional property of Lipschitz functions that
$f_{n}(u)=-\log\Big(\tfrac{1}{n}\sum_{i=1}^{n}r_{n}(x_{i})^{u}\Big)$
is Lipschitz in $u$, with constant only dependent on $c$ and $C$. Thus, we have a uniform Lipschitz bound across all elements of the sequenece $\{f_{n}\}$. It follows that $\{f_{n}\}$ is stochastically equicontinuous on $[0,1]$. ∎
Lemma E.4.
$\{f_{n}\}$ converges uniformly in probability for $u\in[0,1]$
Proof.
This follows from Theorem 1 from Newey (1991) as well as pointwise convergence in Lemma E.2 and stochastic equicontinuity in Lemma E.3. ∎
Finally, using these lemmas, we can complete the proof of our theorem.
Theorem E.2 (Restatement of Theorem 2).
Let $P_{0}$ and $P_{1}$ be distributions with common support and bounded density ratio. Then the Chernoff Information estimator, constructed using a consistent density ratio estimator, is itself consistent.
Proof.
This result follows from Theorem E.1. Continuity and the uniqueness of the maximizer are given by Nielsen (2022). Compactness is given by Lemma E.1. Uniform Convergence in probability is given by Lemma E.4. Hence,
$\widehat{\text{CI}}\rightarrow\text{CI}.$ (60)
Therefore, we have consistency of the Chernoff Information estimator. ∎
Appendix F Experiment Details
F.1 Estimation
To implement the density ratio estimation from Choi et al. (2022), we utilize the codebase provided by the authors  [^5] https://github.com/ermongroup/dre-infinity. Following the success in their paper, we choose to implement the “Joint” model where the model, in addition to time scores, aims to recover data scores. For our interpolation technique, we utilize $p_{\lambda}(x)=\lambda P_{0}(x)+\sqrt{1-\lambda^{2}}P_{1}(x)$. For the model architecture, we utilize the Joint architecture proposed in Appendix F of Choi et al. (2022) which leverages multilayer perceptrons with Exponential Linear Unit (ELU) activations as the modules:
  1. 1.
    Joint (Shared):
    $\mathrm{Linear}(3,256)\;\to\;\mathrm{ELU}\;\to\;\mathrm{Linear}(256,512)\;\to\;\mathrm{chunk}(2)$
    1. (a)
      Time Module:
      $\mathrm{Linear}(256,256)\;\to\;\mathrm{ELU}\;\to\;\mathrm{Linear}(256,256)\;\to\;\mathrm{ELU}\;\to\;\mathrm{Linear}(256,1)$
    2. (b)
      Data Module:
      $\mathrm{Linear}(256,256)\;\to\;\mathrm{ELU}\;\to\;\mathrm{Linear}(256,256)\;\to\;\mathrm{ELU}\;\to\;\mathrm{Linear}(256,2).$
We utilize a Cosine Annealing learning rate scheduler with an initial learning rate of 1e-5 and a batch size of 128. We train all models for 30,000 steps with no importance weighting. All other parameters are standard to the original implementation. All models are trained using NVIDIA L40S GPUs. We highlight an ablation study over hyperparameters in Section I. For the Monte Carlo method, we utilize all available data; we find that this yields more accurate estimation. To solve the infimum, we utilize Brent’s algorithm, although we highlight that this is not required following Lemma 5 (any convex optimization method would suffice).
F.1.1 Time Complexity
The overall time complexity of our Chernoff Information estimator can be given as:
$T_{\text{total}}=T_{\text{DRE}}+n\cdot C_{\text{int}}+n\cdot C_{\text{MC}}+T_{\text{cvx}}$ (61)
where $n$ is the number of samples used to compute Chenroff Information, $T_{\text{DRE}}$ is time spent on density ratio estimation (DRE) in Line 1, $C_{\text{int}}$ is the time required for one integral operations given in Line 2, $C_{MC}$ is the time required for simple addition/floating point operations required in Line 3, and $T_{\text{cvx}}$ is the time required to solve the convex optimization in Line 4, which is a constant time operation. Hence, for other than $T_{\text{DRE}}$, the rest of the operation is done in $O(n)$ and training a neural network for DRE is often the most time-consuming part, which took approximately 30 minutes on a single NVIDIA L40S GPU.
F.2 Gaussian
For all Gaussian experiments, we estimate Chernoff information using the parameters and algorithm specified in F.1. We work with 10,000 samples to perform our estimation. First, to compare to closed form Chernoff Information, we create 2D Gaussians $\mathcal{N}(\textbf{0},\sigma^{2}\mathbf{I})$ and $\mathcal{N}(\textbf{1},\sigma^{2}\mathbf{I})$ and vary the parameter $\sigma^{2}$. For each $\sigma^{2}$, we perform 5 estimates to create error bars.
For more complex 5D Gaussian distributions, we compare to the algorithm provided by Nielsen (2022). We implement the algorithm directly from the paper and use a stopping criteria of 1e-7. We compute Chernoff Information across different levels of noise added to $\mathcal{N}(\textbf{0},\frac{1}{2}\mathbf{I})$ and $\mathcal{N}(\textbf{1},\textbf{I})$. We repeat each trial 5 times to obtain error bars.
For the final experiment, we repeat the previous example; however instead of adding noise, we scale up the dimension. We again compare to Gaussian estimation algorithm for distributions $\mathcal{N}(\textbf{0},\frac{1}{2}\mathbf{I})$ and $\mathcal{N}(\textbf{1},\textbf{I})$ and repeat each trial 5 times to obtain error bars.
F.3 Datasets
Adult.
For the Adult dataset, we select the continuous features, “age", “fnlwgt", “education-num", “capital-gain", “capital-loss", “hours-per-week", for our experiments. We apply a standard scaler, transforming the distributions so that they are centered at 0 with standard deviation of 1. We consider the standard income classification task and use Sex as the sensitive attribute. To obtain Chernoff Information estimates, we train our model as specified in Section F.1. We train each estimate 5 times to create error bars. We compute Chernoff Difference from these Chernoff Information values.
For fairness-accuracy curves, we train split logistic regression models using an 80/20 train-test split for each model. We fix the model for the dominating group (Male) and sweep class weights for the other group, dropping non-Pareto optimal points. We find that the False Negative Rate disparity dominates and thus utilize that in our plot.
HSLS.
Following Jeong et al. (2022), we select a subset of features from the HSLS dataset. We further restrict to continuous features, giving: “X1MTHID",“X1MTHUTI", “X1MTHEFF", “X1FAMINCOME", “X1SCHOOLBEL". We apply a standard scaler, giving distributions centered at 0 with standard deviation of 1. For our sensitive attribute create groups using the Race attribute, giving Asian/White and URM (Other). We utilize the “X1TXMSCR" column to create a binary classification task (top 50th percentile or not). For Chernoff Information and fairness-accuracy curves, we replicate the experiments we conduct on the Adult dataset. We find that the False Positive Rate disparity dominates and thus utilize that in our plot.
MNIST
We replicate the previous experiments using 2D embeddings of MNIST data which were obtained via Principal Component Analysis. For $P_{0}$, we select the digit 0, $P_{1}$, we select the digit $2$, $Q_{0}$, we select the digit $3$ and $Q_{1}$ we select the digit 7.
Dataset Licenses
All datasets used in this work (Adult, HSLS, MNIST) are publicly available and approved for non-commercial research use under their respective data use agreements..
F.4 Mixture of Gaussian
For mixture of Gaussian experiments, we design the distributions as follows. We work with 10,000 samples from each group for our experiments. We utilize the Chernoff Information estimator with the same parameters from Section F.1 and obtain the fairness accuracy curves by replicating the experimental procedure from Section 4.
Mixture 1
$P_{0}$ $=\tfrac{1}{2}\,\mathcal{N}\!\left(\begin{bmatrix}10.5\\ 10.5\end{bmatrix},6.56\,\textbf{I}\right)+\tfrac{1}{2}\,\mathcal{N}\!\left(\begin{bmatrix}10.8\\ 11.2\end{bmatrix},6.43\,\textbf{I}\right),$ (62)
[…]
$Q_{0}$ $=\tfrac{1}{2}\,\mathcal{N}\!\left(\begin{bmatrix}0.5\\ 0.5\end{bmatrix},0.55\,\textbf{I}\right)+\tfrac{1}{2}\,\mathcal{N}\!\left(\begin{bmatrix}0.5\\ 0.5\end{bmatrix},0.45\,\textbf{I}\right),$ (64)
$Q_{1}$ $=\tfrac{1}{2}\,\mathcal{N}\!\left(\begin{bmatrix}3.8\\ 3.8\end{bmatrix},\begin{bmatrix}0.2&0.1\\ 0.1&0.2\end{bmatrix}\right)+\tfrac{1}{2}\,\mathcal{N}\!\left(\begin{bmatrix}2.8\\ 2.8\end{bmatrix},0.55\,\textbf{I}\right).$ (65)
(a) Group $P$
(b) Group $Q$
Figure F.1: (Mixture 1)
Mixture 2
[…]
$P_{1}$ $=\tfrac{1}{2}\,\mathcal{N}\!\left(\begin{bmatrix}3.1\\ 2.4\end{bmatrix},\begin{bmatrix}0.22&0.19\\ 0.19&0.22\end{bmatrix}\right)+\tfrac{1}{2}\,\mathcal{N}\!\left(\begin{bmatrix}2.9\\ 3.2\end{bmatrix},0.24\,\textbf{I}\right),$ (67)
[…]
$Q_{1}$ $=\tfrac{1}{2}\,\mathcal{N}\!\left(\begin{bmatrix}1.3\\ 1.3\end{bmatrix},\begin{bmatrix}9&4\\ 4&9\end{bmatrix}\right)+\tfrac{1}{2}\,\mathcal{N}\!\left(\begin{bmatrix}1.3\\ 1.3\end{bmatrix},8\,\textbf{I}\right).$ (69)
(a) Group $P$
(b) Group $Q$
Figure F.2: (Mixture 2)
Appendix G HSLS
Figure G.1: (HSLS Experiments) (a) Chernoff Information for White/Asian and URM groups decrease as noise ($\eta^{2}$) increases. (b) Chernoff Difference values are larger with a small increase but still remain flat (Case 1). (c) Following the low Chernoff Difference values, are steeper and grow slightly more steep as noise is added.
HSLS.
We examine the behavior of Chernoff Information on the HSLS dataset (Ingels et al., 2011). Following Jeong et al. (2022), we select a subset of the features (further filtering to include only continuous features) to perform a binary classification task (top 50th percentile mathematics test score). This gives a 5 dimensional dataset on which we perform our analysis. For the binary groups, we split the data into two groups based on Race, Asian/White and Under Represented Minorities (URM). We follow a similar approach to the Adult dataset to create the Chernoff Difference plots. Additionally, we follow a similar approach to generate fairness-accuracy plots, however, we find that the false positive rate dominates the error and the Chernoff Exponent, using that as our fairness metric.
Again, the Chernoff Information decays as noise is added, however as opposed to the Adult dataset, there is a much larger disparity between the value across groups. This is reflected as we see much larger values of Chernoff Difference in Figure 1(b) and steeper fairness accuracy curves (Figure 1(c)).
Similar to the Adult dataset, we observe the Chernoff Difference remains relatively stable as the variance of noise ($\eta^{2}$) increases, however, we do note a small increase before the Chernoff Difference begins to slightly decrease (indicative of Case 1). We see that this trend is somewhat reflected, by a very subtle steepening of the fairness accuracy curve (we point the reader to Appendix H for the log-fairness accuracy curve where this very subtle trend can be more easily examined). These observations could resemble Case 1, however, we point out that this trend is very subtle and for the most part, the slopes remain flat. Overall, these experiments suggest that in this dataset, privacy and fairness may be slightly less compatible, although the effect is minor.
Appendix H Supplemental Real Data Figures
Figure H.1: (Adult Dataset) (a) Fairness-Accuracy Curve. (b) Log Fairness-Accuracy Curve. We find that the curves remain very stable, reflected by the small changes in Chernoff Difference.
Figure H.2: (HSLS Dataset) (a) Fairness-Accuracy Curve. (b) Log Fairness-Accuracy Curve. We observe a very slight steepening of the fairness accuracy curves, further emphasized by the log-fairness accuracy curves.
Appendix I Ablation Study
All experiments are performed on 5D Gaussians $\mathcal{N}(\textbf{0},\frac{1}{2}\mathbf{I})$ and $\mathcal{N}(\textbf{1},\mathbf{I})$, with the standard setup mentioned in Section F.1 unless otherwise stated. All plots are Chernoff Information estimates across steps of the training procedure for the density ratio estimation.
Figure I.1: (Learning Rate): Comparison of Chernoff Information Estimation for different learning rates. We find learning rate is the most sensitive hyperparameter and may converge to an incorrect value if too small or be unstable if too large. We explore a selection of learning rates and find that 1e-5 works well under Gaussian and tabular settings.
Figure I.2: (Batch Size) Comparison of Chernoff Information Estimation for different batch sizes. We find that batch size has a more subtle effect on the estimation outcome. Larger batch sizes tend to be slightly more beneficial for estimation. However, the sample sizes of the tabular datasets we use are smaller than the synthetic Gaussian experiments, causing us to scale the batch size down to 128 for our experiments. These larger batch sizes lead to explosion of the density ratio estimates in tabular settings. We find that this slightly smaller value performs well on both the Gaussian and Tabular settings and prevents the density ratio from exploding in the later case.
Figure I.3: (Sample Size) Comparison of Chernoff Information Estimation for sample sizes. We find that even down to 1000 samples (sample sizes similar to the smallest group in the tabular datasets), the estimation remains fairly accurate. When there are not enough samples, the estimate begins to degrade, as we can begin observing with 500 samples. This sample size and dimension relationship is further highlighted in Figure 3(c).